# The ALPINE-CRISTAL-JWST Survey: Revealing Less Massive Black Holes in High-Redshift Galaxies

### v1

@@ 3.1 Assumptions and Simulations
For the mock spectra, we start from the narrow H line. Given a stellar mass , we estimate the star formation rate (SFR) using the main sequence relation from Speagle et al. (2014, , eq. 28). We assume a redshift of , based on the redshift distribution of the ALPINE-CRISTAL-JWST sample. From the SFR, we derive the H line luminosity using the correlation from Zhuang & Ho (2019, , eq. 1), combining the excitation conversion relation from Kennicutt (1998) with the initial mass function from Kroupa (2001). We adopt fixed line ratios of and , typical values for weak AGNs at high redshift (See Fig. 11). The flux ratio of the [N ii] 6550, 6585 is fixed to its theoretical value of 2.96 (see Section 2.2). While the [S ii] 6718, 6733 ratio varies with electron density, we adopt a value of 1.0. This choice has a negligible impact on our results due to the intrinsic weakness of these lines. The FWHM of the narrow lines is set to 350 km s-1 after including the instrumental resolution, which is the median FWHM of the narrow H line in our sample. The narrow line width is dominated by the spectral resolution, the small variation between different is negligible.
@@ 4.2 Six other AGN Candidates
DC_683613 (HZ3/CRISTAL-05) exhibits a broad H component with FWHMcorr = 678 km s-1 and . The fitting to the [O iii] region rejects an outflow scenario. We note that this target has a relatively strong [N ii] and [S ii] line flux with and , significantly higher than the rest of our sample (see Sec. 5.4). With a relatively lower , this target situates well in the local composite region of the BPT diagram (Baldwin et al., 1981). We note that it is confirmed as a merger system according to its [C ii] observation (Posses et al., 2024) and has an off-centre dust geometry traced by far-infrared observation (Killi et al., 2024). Disfavoring the outflow cause of the dust geometry, this observation further indicates the possibility of AGN feedback in this system.
@@ 5.3 Morphology of AGN host galaxies
The major component exhibits more complex features. Davies et al. (2025, in prep) report evidence for a biconical outflow centred at aperture d, extending in the east-west direction. While we detect broad H emission at position d, this feature fails our line flux ratio criterion (eq. 4) and is likely associated with the outflow. At position b, however, the absence of [O iii] outflow signatures combined with the broad H component suggests an AGN origin. The location of b near the dynamical centre of the galaxy further supports this interpretation, as does its composite classification in the BPT diagram.
@@ 6.2 Gas Kinematics of DC_536534
The line ratios of the two fitted components differ markedly, with for the blue core and for the red component. Despite this substantial difference, both ratios are consistent with the broad range observed in other samples (e.g., Silverman et al., 2009), which allows for the interpretation that the lobe is a genuine astrophysical feature rather than an artifact of our spectral decomposition.
To complement our analysis of ionized gas, we examined the neutral gas distribution through [C ii] observations of DC_536534. We combined data from ALPINE (Béthermin et al., 2020) and CRISTAL (Herrera-Camus et al., 2025) programs to achieve enhanced dynamic range. Following continuum subtraction, we generated spectral cubes using tclean in two modes: cube mode with 20 km/s channels and mfs mode integrating over 2 FWHM from Béthermin et al. (2020). The final [C ii] map, produced with robust Briggs weighting yielding a 0.47″ beam (Briggs, 1995), is presented in the lower-left panel of Fig. 12 (see Liu et al., 2025, in prep, for detailed reduction procedures). The map reveals an additional lobe extending south-eastward from the galaxy centre. We generated velocity-divided contours by integrating (3) from the line centre to the red- and blue-shifted components, shown in magenta (red-shifted) and cyan (blue-shifted). While these contours align with the rotation axis derived from ionized gas, the kinematics of the lobe are enhanced by outflow signatures.
The [C ii] emission can reveal a more extensive neutral gas structure than the ionized gas tracers, with minimal line-of-sight obscuration effects. The [C ii] map of DC_536534 shows a distinct outflow signature in the southeast quadrant of the galaxy, possibly driven by AGN activity. At approximately 1.5 kpc south of the galaxy centre, we detect redshifted component at the base of this outflow in both [O ii] and [O iii] lines. Compared to the central emission, this infalling component is characterized by strong [O ii] line flux but a significantly weaker [O iii] signal. We consider two alternative scenarios for this distinctive structure. First, it could originate from localized star formation triggered by a minor merger. This interaction would induce the general outflow seen in [C ii] and simultaneously trigger star formation in the local, interactive region, explaining the low-ionization [O ii] emission. Alternatively, this infalling gas may result from the AGN-driven outflow interacting with the interstellar medium, potentially providing a gas recycling mechanism to sustain future AGN activity.
@@ 6.3 Stellar Mass Estimation
The stellar mass estimation used in this work is derived through SED fitting with pre-JWST broad-band photometry from Faisst et al. (2020). These spatially integrated measurements can be affected by strong emission lines in actively star-forming galaxies, potentially bias the stellar mass estimation. To explore this issue, we analysed the relations using spatially resolved stellar mass measurements from Tsujita et al. (2025, in prep). Their analysis employs Prospector to simultaneously fit NIRCam photometry and NIRSpec IFU spectroscopy at 0.6 kpc resolution. For each AGN candidate, we summed the stellar masses of constituent pixels, considering only pixels within the host galaxy for merger systems and corrected the flux loss due to the mask of low S/N pixels.
As shown in Fig. 13, for our AGN candidate sample, these spatially resolved estimates are systematically lower by 0.6 dex compared to the integrated broad-band measurements. It is worth noting that the general deviation for the 18 ALPINE galaxies is only 0.35 in average, indicating extended difference in our AGN candidates.
@@
Additional uncertainty could arise from strong narrow-line emission. Another spatially resolved SED fitting result based solely on the broad-band images (Li et al., 2024b) finds no offsets between the derived from the integrated measurements regardless the different SED models. Further discussion about the origin of these differences are beyond the scope of this study, which will be further addressed in Tsujita et al. (2025, in prep).
@@ Acknowledgements
The data underlying this article are available in the article.

### v2

@@ 3.1 Assumptions and Simulations
For the mock spectra, we start from the narrow H line. Given a stellar mass , we estimate the SFR using the main sequence relation from Speagle et al. (2014, , eq. 28). We assume a redshift of , based on the redshift distribution of the ALPINE-CRISTAL-JWST sample. From the SFR, we derive the H line luminosity using the correlation from Zhuang & Ho (2019, , eq. 1), combining the excitation conversion relation from Kennicutt (1998) with the initial mass function from Kroupa (2001). We adopt fixed line ratios of and , typical values for weak AGNs at high redshift (See Fig. 11). The flux ratio of the [N ii] 6550, 6585 is fixed to its theoretical value of 2.96 (see Section 2.2). While the [S ii] 6718, 6733 ratio varies with electron density, we adopt a value of 1.0. This choice has a negligible impact on our results due to the intrinsic weakness of these lines. The FWHM of the narrow lines is set to 350 km s-1 after including the instrumental resolution, which is the median FWHM of the narrow H line in our sample. The narrow line width is dominated by the spectral resolution, the small variation between different is negligible.
@@ 4.2 Six other AGN Candidates
DC_683613 (HZ3/CRISTAL-05) exhibits a broad H component with FWHMcorr = 678 km s-1 and . The fitting to the [O iii] region rejects an outflow scenario. We note that this target has a relatively strong [N ii] and [S ii] line flux with and , significantly higher than the rest of our sample (see Sec. 5.4). With a relatively lower , this target situates in the local composite region of the BPT diagram (Baldwin et al., 1981). We note that it is confirmed as a merger system according to its [C ii] observation (Posses et al., 2024) and has an off-centre dust geometry traced by far-infrared observation (Killi et al., 2024). Disfavoring the outflow cause of the dust geometry, this observation further indicates the possibility of AGN feedback in this system.
@@ 5.3 Morphology of AGN host galaxies
The major component exhibits more complex features. Davies et al. (2025, in prep) report evidence for a biconical outflow centred at aperture d, extending in the east-west direction. While we detect broad H emission at position d, this feature fails our line flux ratio criterion (eq. 4) and is likely associated with the outflow. At position b, however, the absence of [O iii] outflow signatures combined with the broad H component suggests an AGN origin. The location of b near the dynamical centre of the galaxy further supports this interpretation.
@@ 6.2 Gas Kinematics of DC_536534
The line ratios of the two fitted components differ markedly, with for the blue core and for the redshifted component. Despite this substantial difference, both ratios are consistent with the broad range observed in other samples (e.g., Silverman et al., 2009), which allows for the interpretation that the lobe is a genuine astrophysical feature rather than an artifact of our spectral decomposition.
To complement our analysis of ionized gas, we examined the neutral gas distribution through [C ii] observations of DC_536534. We combined data from ALPINE (Béthermin et al., 2020) and CRISTAL (Herrera-Camus et al., 2025) programs to achieve enhanced dynamic range. Following continuum subtraction, we generated spectral cubes using tclean in two modes: cube mode with 20 km/s channels and mfs mode integrating over 2 FWHM from Béthermin et al. (2020). The final [C ii] map, produced with robust Briggs weighting yielding a 0.47″ beam (Briggs, 1995), is presented in the lower-left panel of Fig. 12 (see Liu et al., 2025, for detailed reduction procedures). The map reveals an additional lobe extending south-eastward from the galaxy centre. We generated velocity-divided contours by integrating (3) from the line centre to the red- and blue-shifted components, shown in magenta (redshifted) and cyan (blueshifted). While these contours align with the rotation axis derived from ionized gas, the kinematics of the lobe are enhanced by blueshifted signatures.
The [C ii] emission of DC_536534, which traces extensive neutral gas structures with minimal obscuration, reveals a distinct lobe extending southeast from the galaxy center. The kinematics of this feature are broadly aligned with the rotation, providing evidence that the structure is a tidal feature from an ongoing minor merger (Villalobos & Helmi, 2008). At the base of this structure, approximately 1.5 kpc south of the center, we detect a redshifted component in both [O ii] and [O iii] lines. This component is characterized by strong [O ii] flux but a significantly weaker [O iii] signal. The discrepancy between the bulk gas motion (traced by [C ii]) and the ionized gas (traced by [O ii]) can be naturally explained by a moderate-velocity shock front, which heated the interstellar medium and result in the intense, low-ionization [O ii] emission and its unique velocity signature (Rich et al., 2011). An alternative, or perhaps complementary, scenario is that the interaction has triggered a localized burst of star formation in this region, as evident by the low-ionization [O ii] emission.
@@ 6.3 Stellar Mass Estimation
The stellar mass estimation used in this work is derived through SED fitting with pre-JWST broad-band photometry from Faisst et al. (2020). These spatially integrated measurements can be affected by strong emission lines in actively star-forming galaxies, potentially bias the stellar mass estimation. To explore this issue, we analysed the relations using spatially resolved stellar mass measurements from Tsujita et al. (2025, submitted). Their analysis employs Prospector to simultaneously fit NIRCam photometry and NIRSpec IFU spectroscopy at 0.6 kpc resolution. For each AGN candidate, we summed the stellar masses of constituent pixels, considering only pixels within the host galaxy for merger systems and corrected the flux loss due to the mask of low S/N pixels.
As shown in Fig. 13, for our AGN candidate sample, these spatially resolved estimates are systematically lower by 0.6 dex compared to the integrated broad-band measurements. It is worth noting that the spatially resolved stellar masses for the full ALPINE sample are also presented in Tsujita et al. (2025, submitted). The general deviation for the 18 ALPINE galaxies is only 0.35 dex on average. This indicates a more significant discrepancy in our AGN candidates compared to the general galaxy population at similar redshifts.
@@
Additional uncertainty could arise from strong narrow-line emission. Another spatially resolved SED fitting result based solely on the broad-band images (Li et al., 2024b) finds no offsets between the derived from the integrated measurements regardless the different SED models. Further discussion about the origin of these differences are beyond the scope of this study, which will be further addressed in Tsujita et al. (2025, submitted).
@@ Acknowledgements
The data underlying this research are available in the article.